WAC Notes February 3 2020

From PREX Wiki
Jump to: navigation, search
Documentation
HOW TOs for shift crew
Expert Tools
All Expert Contacts



PREX Main << Weekly Analysis Coordinator << WAC Notes CREX << Instructions

WAC Notes February 2 2020 << >> WAC Notes February 4 2020

February 3 2020

Daily Meeting BlueJeans


Production Run List - Slug 127 - IHWP OUT - Flip-RIGHT - Production

continue slug 127 from February 2:

Production Run List - Slug 128 - IHWP IN - Flip-RIGHT - Production

start this slug at about 6 am

  • Good - Run 6163
  • Good - Run 6164
  • Good - Run 6165
  • Good - Run 6166
  • Good - Run 6168
  • Good - Run 6176
  • Good - Run 6177 - Nonzero analog BCM DD mean value. There is clear trending up of DD value. will cut some trip recoveries
    • Later: Adding Cuts for some trips cuts are needed to clean up the DD. Events 0:65k, 100k:140k, 225k:240k, 270k:280k
  • Good - Run 6178
  • Good - Run 6179 - Similar problems as 6177, but it seems harder to tell where the problem is. There is uncut beam motion though, so adding that now.
    • Later: Adding Cuts Adding bad events cut for position issues at 150:200k and that's it.

continue slug 128 on February 4:

Non-Production Runs and Bad Runs

Pedestal/Calibrations Updates

Issues

  • double peak in bpm X appears since run 6161, continues to most recent run
  • raw asym of url goes up to huge value
  • 3 (12:00 - 15:00) hours of hall C spin dance
  • LQ2 tripped off at 19:07
  • it looks like the bpm X double peak just gone after we return to production run 6176, though we still don't know why there comes the double peak and how it goes away?
  • for run 6176-6179, MCC can't deliver beam larger than 120 uA, so we just run at 120 uA

Updates